If Lane invested $22,000, part at 8% and part at 7%. If the total interest at the end of the year is $1,710, how much did she invest at 8%
Answer by checkley71(8403)

.08X+.07(22000-X)=1710
.08X+1540-.07X=1710
.01X=1710-1540
.01X=170
X=170/.01
X=17000 AMOUNT INVESTED @ 8%
22000-17000=5000 INVESTED @ 7%
PROOF
.08*17000+.07*5000=1710
1360+350=1710
1710=1710